• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name folts

    Log entry time 18:08:19 on April 18, 2007

    Entry number 199512

    keyword=late entry/ tech

    Earlier today we changed the target camera (nice job Scot)

    Moved the left spectrometer to forward of 120 degrees and installed the pin.

    Moved the right spectrometer to larger than 15 degrees and installed the 14 degree pin. during the rightspectrometer move I had to depower and repower the bogies and reboot the bogie crate to get the page to come up. FYI

    I did not lower the lead flow to right Q1 do to time restrictions.

    A copy of this log entry has been emailed to: fansler@jlab.org